starting from maxwell's equations derive wave equation for a uniform plane wave travelling in free space.

1
Expert's answer
2022-03-22T15:48:20-0400

"\nabla E=-\frac{\rho}{\epsilon_0}"

"E=E_0e^{kx-wt}"


"=\nabla E=\frac{d(E_0e^{kx-wt})}{dx}=kE_0e^{kx-wt}=kE"

"kE=-\frac{\rho}{\epsilon_0}\\E=-\frac{\rho}{k\epsilon_0}"

Wave equation